OpenStudy (anonymous):

4. Suppose y varies directly with x, and y=24 when x=8. What is the value of y when x=10? (1 point) A. 10 B. 30 C. 80 D. 240 5. What is an equation for the line with the slope 2/3 and y-intercept 9? (1 point) A. y= 2/3x B. y= 9x C. y= 2/3x +9 D. y= 9x +2/3 6. What is an equation in slope-intercept form for the line that passes through the points (1, -3) and (3, 1)? (1 point) A. y=3x +1 B. y= x- 3 C. y= 2x +5 D. y= 2x -5

OpenStudy (dmndlife24):

y = kx

OpenStudy (dmndlife24):

24 = k8 k = 3 y = (8)10 y = 80

OpenStudy (jhannybean):

4. direct variation: y = kx . K is the constant of variation, so solving for k, we get, k = y/x. You're given y = 24 and x = 8, plugging these in: k = 24/8 = 3 Now we're asked for y = kx when x= 10 y = 3(10) = ?

OpenStudy (jhannybean):

@dmndlife24 you solved for k and got 3 but plugged in 8 for k.

OpenStudy (dmndlife24):

2. y = mx +b